Find the derivative of the function. f(x) = (ln(x + 5)) f(c) = PreviewFind the derivative of the function. f(t) = ť(In(t))? f(t) = PreviewIf f(a) = 8 ln(4x), find a. f b. Rounded to the nearest whole number: f(e) c. Rounded to the nearest whole number: f(e) = d.
Answer #1

if f(x) = 8 In (40) a) !64) - b.) fle) - 816 (4e) Le) = 19 c) fle) = 8x42 8. 42 de a 8x4 ye fe) = 3 d.) equation of tangent y> if f(x) = 8 11 (4x) f ) (M (+5)) f(x) 4 (In (ats da x+5 41n(x+5 Hfix = 4/10 (2+6) © flt) = + (In(+3)] $19) = 2+(\n (t?))